# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ 6 Data Analysis
Table 5: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,...,21$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$)
@@
$\beta_{1}$ 5.743 0.784 (4.298, 7.330) 6.211 1.268 (4.179, 8.928)
$\beta_{2}$ 5.649 0.768 (4.255, 7.152) 6.018 1.279 (3.897, 8.845)
$\beta_{3}$ 5.724 0.758 (4.269, 7.289) 6.143 1.141 (4.360, 8.676)
$\beta_{4}$ 5.692 0.780 (4.237, 7.328) 6.088 1.142 (4.203, 8.632)
$\beta_{5}$ 5.671 0.774 (4.146, 7.310) 6.019 1.197 (4.047, 8.581)
$\beta_{6}$ 5.633 0.761 (4.225, 7.219) 5.953 1.154 (3.964, 8.361)
$\beta_{7}$ 5.584 0.754 (4.088, 7.015) 5.924 1.149 (3.989, 8.442)
$\beta_{8}$ 5.561 0.763 (4.075, 7.007) 5.809 1.171 (3.870, 8.409)
$\beta_{9}$ 5.579 0.763 (4.048, 7.118) 5.873 1.113 (3.980, 8.048)
$\beta_{10}$ 5.528 0.743 (4.027, 6.960) 5.770 1.108 (3.910, 8.085)
$\beta_{11}$ 5.415 0.761 (3.930, 6.934) 5.665 1.100 (3.791, 7.968)
$\beta_{12}$ 5.402 0.742 (3.907, 6.872) 5.630 1.147 (3.681, 7.926)
$\beta_{13}$ 5.252 0.749 (3.718, 6.727) 5.404 1.149 (3.467, 7.879)
$\beta_{14}$ 5.212 0.744 (3.698, 6.693) 5.316 1.153 (3.203, 7.660)
$\beta_{15}$ 5.244 0.729 (3.741, 6.537) 5.402 1.149 (3.430, 7.872)
$\beta_{16}$ 5.105 0.745 (3.564, 6.477) 5.194 1.202 (3.099, 7.615)
$\beta_{17}$ 5.051 0.750 (3.632, 6.429) 5.099 1.204 (2.986, 7.531)
$\beta_{18}$ 4.987 0.783 (3.364, 6.405) 5.033 1.179 (2.916, 7.587)
$\beta_{19}$ 4.915 0.800 (3.171, 6.354) 4.942 1.212 ( 2.739, 7.298)
$\beta_{20}$ 4.907 0.777 (3.321, 6.329) 4.870 1.213 (2.661, 7.278)
$\beta_{21}$ 4.853 0.780 (3.208, 6.329) 4.854 1.235 (2.586,7.146)
$\alpha$ 0.844 0.028 (0.787, 0.901) 0.832 0.049 (0.716, 0.899)
$\sigma_{\epsilon}^{2}$ 0.113 0.017 (0.079, 0.149) 0.127 0.048 (0.081, 0.225)
Figure 8: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by Semi-parametric method.
Figure 9: For each row, the left side plot is for histogram for Residual lifetimes, middle one is for trace plot of generated samples and right side denotes autocorrelation plot produced by parametric method.
@@
The mean, standard deviation (s.d) and 95% credible interval (C.I) of posterior samples of random effects for each of 21 units, fixed effect $\alpha$ and variance of measurement error $\sigma_{\epsilon}^{2}$ are presented in Table 5. Observe that for some units, posterior mean of random effect is significantly higher than some other units. For example for unit 1, the estimated posterior mean for random effects is much higher than that of unit 20, which can motivate us to analyze this data using the proposed semi-parametric model.
@@
Table 6: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method
@@
Unit True value Parametric method Semi-parametric method

### v2

@@ 6.2 Case 2
Table 9: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 2 for both parametric and semi-parametric method
@@
$\beta_{1}$ 7.220 0.350 (6.514, 7.907) 7.196 0.325 (6.595, 7.887)
$\beta_{2}$ 3.760 0.339 (3.105, 4.415) 3.732 0.339 (3.067, 4.456)
$\beta_{3}$ 7.157 0.329 (6.518, 7.791) 7.165 0.335 (6.544, 7.862)
$\beta_{4}$ 2.020 0.341 (1.374, 2.710) 2.047 0.348 (1.394, 2.701)
$\beta_{5}$ 6.157 0.334 (5.514, 6.832) 6.148 0.346 (5.474, 6.832)
$\beta_{6}$ 2.404 0.339 (1.794, 3.090) 2.373 0.349 (1.710, 3.089)
$\beta_{7}$ 6.943 0.337 (6.266, 7.582) 6.950 0.329 (6.361, 7.608)
$\beta_{8}$ 2.016 0.335 (1.361, 2.660) 2.024 0.342 (1.370, 2.666)
$\beta_{9}$ 1.927 0.325 (1.313, 2.581) 1.917 0.335 (1.249, 2.609)
$\beta_{10}$ 2.026 0.342 (1.357, 2.733) 2.002 0.343 (1.348, 2.709)
$\alpha$ 2.062 0.105 (1.857, 2.259) 2.069 0.107 (1.845, 2.265)
$\sigma_{\epsilon}^{2}$ 0.989 0.113 (0.792, 1.225) 0.983 0.116 (0.783, 1.252)
@@
Table 10: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 3 for parametric and prior 5 for semi-parametric method
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 7.237 0.337 (6.543, 7.888) 7.236 0.338 (6.655, 7.952)
$\beta_{2}$ 3.745 0.340 (3.075, 4.378) 3.749 0.344 (3.062, 4.450)
$\beta_{3}$ 7.164 0.351 (6.539, 7.848) 7.172 0.349 (6.460, 7.846)
$\beta_{4}$ 2.026 0.355 (1.338, 2.754) 2.033 0.342 (1.382, 2.729)
$\beta_{5}$ 6.171 0.342 (5.528, 6.838) 6.152 0.342 (5.499, 6.806)
$\beta_{6}$ 2.390 0.342 (1.735, 3.075) 2.389 0.355 (1.675, 3.127)
$\beta_{7}$ 6.979 0.336 (6.303, 7.644) 6.975 0.346 (6.318, 7.659)
$\beta_{8}$ 1.995 0.325 (1.345, 2.628) 2.007 0.352 (1.318, 2.742)
$\beta_{9}$ 1.906 0.352 (1.188, 2.551) 1.906 0.337 (1.255, 2.558)
$\beta_{10}$ 2.009 0.334 (1.371, 2.663) 2.017 0.347 (1.359, 2.708)
$\alpha$ 2.065 0.106 (1.851, 2.268) 2.064 0.110 (1.833, 2.259)
$\sigma_{\epsilon}^{2}$ 0.992 0.112 (0.802, 1.236) 0.988 0.119 (0.772, 1.242)
It is evident from Tables 9 and 10, that 95% credible interval produced by both parametric and proposed semi-parametric do not contain true value for error variance under different prior combinations. On the other hand 95% credible intervals for both the method contains the true value of the fixed effect parameter. We observe that for each unit, the posterior mean of random effects is very close for both the methods. The posterior estimates are not significantly influenced by prior distributions as the number of observations per unit increases. We consider prior 2 to compute residual lifetime distribution function for both the model.
@@
Table 11: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method with $m=31$
@@
. Unit True value Parametric Semi-parametric median 95% Predictive interval median 95% Predictive interval 1 0.194 0.378 (0.100, 0.643) 0.381 (0.109, 0.678) 2 1.095 1.657 (0.984, 2.289) 1.663 (1.026, 2.349) 3 0.188 0.391 (0.099, 0.657) 0.391 (0.125, 0.677) 4 2.987 3.970 (2.350, 6.543) 3.846 (2.134, 6.024) 5 0.286 0.630 (0.315, 1.023) 0.606 (0.252, 0.934) 6 2.556 3.174 (1.832, 4.619) 3.189 (1.730, 4.488) 7 0.171 0.433 (0.159, 0.759) 0.432 (0.172, 0.716) 8 2.862 3.965 (2.225, 6.019) 3.950 (2.359, 5.796) 9 2.717 4.183 (2.439, 6.174) 4.171 (2.394, 6.284) 10 3.554 3.871 (2.205, 6.123) 4.063 (2.347, 6.236)
Table 12: KS distance between true and approximated residual lifetime distribution for Bayesian parametric method and proposed Bayesian semi-parametric method
. Unit $m=11$ $m=31$ Parametric Semi-parametric Parametric Semi-parametric 1 0.166 0.810 0.291 0.283 2 0.997 0.622 0.135 0.160 3 0.980 0.720 0.098 0.122 4 1.000 0.661 0.227 0.284 5 1.000 0.577 0.294 0.254 6 0.989 0.841 0.487 0.473 7 0.334 0.607 0.154 0.163 8 1.000 0.677 0.219 0.217 9 0.325 0.551 0.244 0.266 10 0.999 0.827 0.532 0.463
Table 13: Prediction accuracy for Parametric and Semi-Parametric method
. Scenario $M_{1}$ $M_{2}$ Parametric Semi-parametric Parametric Semi-parametric RMSE MAE RMSE MAE RMSE MAE RMSE MAE $n=10,m=11$ 2.258 1.306 0.477 0.572 2.202 1.299 0.497 0.582 $n=10,m=31$ 0.734 0.777 0.726 0.780 0.729 0.779 0.749 0.792 $n=30,m=11$ 0.865 0.793 0.951 0.837 0.872 0.797 0.941 0.837 $n=30,m=31$ 0.872 0.794 0.879 0.799 0.871 0.798 0.888 0.803 $n=50,m=11$ 0.925 0.791 1.131 0.884 0.934 0.796 1.123 0.883 $n=50,m=31$ 0.858 0.822 0.846 0.821 0.852 0.821 0.856 0.824
Predictive intervals are constructed using the 95% highest probability density region of the generated samples from residual lifetime distribution for individual unit. The estimates for residual lifetime along with the true residual lifetime values for each unit are presented in Table 11. The histograms of the generated residual lifetimes for units 1-3 along with the trace plot of and autocorrelation plot of generated samples for semi-parametric and parametric method are provided in Figure 5-6, given in the Appendix A to this article. It is evident from Table 11 that for most of the units both parametric and semi-parametric methods produce intervals containing the true value. Predictive interval produced by parametric method fails to contain true value for unit 5 whereas the semi-parametric method fails to contain true value for unit 7. We find that both parametric and semi-parametric method fails to accurately estimate the tail of the true distribution in both case 1 and 2, due to the reason that the each of the 3-component mixture in both the cases are skewed, but the methods relies on either mixture of normals or unimodal normal distributions.
Kolmogorov-Smirnov (KS) distance obtained for $m=11,31$ and $n=10$ is given in Table 12. We observe that for the slowly degrading units for example unit 4, 5 KS statistic is very large for parametric method compared to the semi-parametric method when $m=10$. The reason for this is that the mixture of normal is able to model the heterogeneous nature of the true distribution, whereas the unimodal distribution for parametric method is unable to estimate this when there are not enough observations per unit. On the other side, KS distance for parametric method is lesser for faster degrading units for $m=10$. The same pattern also evident for Case 1, which indicates that the parametric method performs well for the units coming from mixing component with highest mean. As number of observations per individual increases value of KS statistic decreases for both the method. To assess the overall prediction accuracy, RMSE and MAE are calculated for each combination of $n$ and $m$. Similar to Case 1, It is evident from Table 13 that, as the number of observations per unit increases for $n=30,50$, both RMSE and MAE decreases under $M_{1}$ and $M_{2}$. Both the proposed semi-parametric and parametric method for $M_{1}$ and $M_{2}$, is nearly equal as both the observations and sample size increases. For both case 1 and 2 method $M_{2}$ can produce non-negative values which can decrease the overall RMSE or MAE, but method $M_{1}$ will produce positive values which is appropriate for residual life according to our assumption. In both scenarios for case 1 and 2, it is observed that the estimated median of residual lifetime is higher than true value. The reason for this is the residual lifetime distribution functions derived to simulate observations involves terms which are average of many distribution functions, as a result this induces a much longer upper tail and hence higher values are simulated.
6.3 Case 3
We consider a 2-component mixture of normal distribution for random effect for Case 3, where mean of first component is 3 and the second component is 6. Degradation observation generated for this case is given in Figure 1(c). The mean, standard deviation (s.d) and 95% Credible interval of random effects for each of 10 units, fixed effect $\alpha$ and variance of error $\sigma_{\epsilon}^{2}$ with respect to the different prior combinations for both the parametric and semi-parametric model are given in Tables 14 and 15. The trace plot and autocorrelation plot for the generated mcmc samples of $\alpha$, $\beta_{i}$, $i=1,2,3$ and $\sigma_{\epsilon}^{2}$ for semi-parametric and parametric method are provided in Figure 17 in Appendix B.
Table 14: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 2
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 2.483 0.294 (1.906, 3.088) 2.494 0.296 (1.925, 3.079)
$\beta_{2}$ 6.737 0.324 (6.150, 7.400) 6.752 0.312 (6.125, 7.347)
$\beta_{3}$ 2.841 0.288 (2.277, 3.408) 2.875 0.287 (2.313, 3.421)
$\beta_{4}$ 3.014 0.282 (2.466, 3.575) 3.029 0.293 (2.464, 3.608)
$\beta_{5}$ 3.589 0.282 (3.041, 4.126) 3.604 0.284 (3.055, 4.170)
$\beta_{6}$ 5.635 0.290 (5.100, 6.234) 5.651 0.289 (5.103, 6.219)
$\beta_{7}$ 3.101 0.291 (2.542, 3.690) 3.118 0.280 (2.584, 3.668)
$\beta_{8}$ 2.874 0.281 (2.329, 3.400) 2.900 0.299 (2.342, 3.484)
$\beta_{9}$ 3.165 0.287 (2.634, 3.737) 3.163 0.291 (2.622, 3.747)
$\beta_{10}$ 6.300 0.301 (5.713, 6.871) 6.303 0.301 (5.684, 6.863)
$\alpha$ 2.033 0.090 (1.859, 2.201) 2.026 0.089 (1.853, 2.203)
$\sigma_{\epsilon}^{2}$ 0.728 0.091 (0.572, 0.917) 0.729 0.087 (0.571, 0.911)
Table 15: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 3 for parametric and prior 5 for semi-parametric method
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 2.488 0.293 (1.928, 3.096) 2.463 0.271 (1.964, 3.022)
$\beta_{2}$ 6.752 0.335 (6.117, 7.417) 6.764 0.337 (6.100, 7.417)
$\beta_{3}$ 2.835 0.287 (2.275, 3.438) 2.843 0.297 (2.268, 3.442)
$\beta_{4}$ 2.995 0.293 (2.453, 3.570) 2.996 0.293 (2.420, 3.604)
$\beta_{5}$ 3.588 0.304 (3.014, 4.191) 3.588 0.298 (3.026, 4.171)
$\beta_{6}$ 5.637 0.296 (5.050, 6.214) 5.644 0.281 (5.082, 6.195)
$\beta_{7}$ 3.101 0.290 (2.557, 3.653) 3.091 0.288 (2.537, 3.656)
$\beta_{8}$ 2.883 0.307 (2.295, 3.467) 2.896 0.299 (2.298, 3.467)
$\beta_{9}$ 3.150 0.299 (2.574, 3.768) 3.160 0.290 (2.632, 3.721)
$\beta_{10}$ 6.304 0.292 (5.708, 6.880) 6.307 0.281 (5.776, 6.858)
$\alpha$ 2.034 0.095 (1.837, 2.215) 2.032 0.093 (1.837, 2.201)
$\sigma_{\epsilon}^{2}$ 0.727 0.092 (0.568, 0.924) 0.724 0.091 (0.556, 0.909)
Observe that 95% credible interval produced by both parametric and proposed semi-parametric contain true fixed effect value under each prior combinations. The estimates for fixed effect parameter are closer to true value as the number of observations per unit increases, whereas the posterior mean for error variance seems to higher than true value for both models. We observe that for fast degrading units, the difference between posterior means of random effect for parametric and semi-parametric method decreases as number of observations per unit increases. We consider prior 2 to compute residual lifetime distribution function for both the model.
Table 16: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method with $m=31$
. Unit True value Parametric Semi-parametric median 95% Predictive interval median 95% Predictive interval 1 0.780 1.844 (0.965, 2.866) 1.772 (0.929, 2.766) 2 0.017 0.158 (0.004, 0.367) 0.165 (0.002, 0.379) 3 1.017 1.409 (0.734, 2.267) 1.421 (0.808, 2.276) 4 0.656 1.274 (0.643, 1.978) 1.292 (0.604, 1.987) 5 0.484 0.947 (0.469, 1.533) 0.936 (0.429, 1.463) 6 0.043 0.257 (0.008, 0.517) 0.237 (0.004, 0.500) 7 0.798 1.234 (0.664, 2.038) 1.217 (0.574, 1.918) 8 1.048 1.421 (0.703, 2.161) 1.403 (0.715, 2.180) 9 0.497 1.212 (0.618, 1.904) 1.215 (0.674, 1.955) 10 0.010 0.158 (0.001, 0.361) 0.146 (0.004, 0.367)
Table 17: KS distance between true and approximated residual lifetime distribution for Bayesian parametric method and proposed Bayesian semi-parametric method
. Unit $m=11$ $m=31$ Parametric Semi-parametric Parametric Semi-parametric 1 0.553 0.543 0.632 0.604 2 0.065 0.076 0.116 0.083 3 0.267 0.276 0.188 0.193 4 0.182 0.183 0.170 0.197 5 0.388 0.393 0.177 0.198 6 0.094 0.079 0.156 0.078 7 0.268 0.260 0.388 0.413 8 0.542 0.570 0.271 0.293 9 0.161 0.164 0.177 0.189 10 0.176 0.164 0.167 0.140
Table 18: Prediction accuracy for Parametric and Semi-Parametric method
. Scenario $M_{1}$ $M_{2}$ Parametric Semi-parametric Parametric Semi-parametric RMSE MAE RMSE MAE RMSE MAE RMSE MAE $n=10,m=11$ 0.474 0.626 0.470 0.621 0.474 0.620 0.462 0.612 $n=10,m=31$ 0.530 0.676 0.515 0.667 0.536 0.675 0.523 0.666 $n=30,m=11$ 0.696 0.796 0.708 0.803 0.689 0.794 0.701 0.801 $n=30,m=31$ 0.559 0.721 0.633 0.771 0.556 0.718 0.633 0.771 $n=50,m=11$ 0.561 0.692 0.824 0.871 0.559 0.688 0.820 0.869 $n=50,m=31$ 0.500 0.678 0.726 0.830 0.499 0.675 0.726 0.829
Predictive intervals based on the 95% highest probability density region of the generated samples for the individual units are presented in Table 16. The histograms of the generated residual lifetimes for units 1-3 along with the trace plot of and autocorrelation plot of generated samples for semi-parametric and parametric method for this case are provided in Figure 7-8 respectively in the Appendix A to this article. For most of the units, the produced predictive intervals for both parametric and semi-parametric methods contain the true value. Predictive interval produced by both parametric and semi-parametric method fails to contain true value for unit 1 and 9.
Kolmogorov-Smirnov (KS) statistic obtained for $m=11,31$ and $n=10$ is given in Table 17. For most of the slowly degrading units for example unit 3, 8 value of KS statistic decreases for both the method as number of observations per unit increases. The same also happens for the fast degrading units. So, in this case performance of both the methods is quite similar, because the heterogeneity is comparatively less for this case. RMSE and MAE are calculated for each combination of $n$ and $m$. Similar to Case 1, It is evident from Table 13 that, as the number of observations per unit increases for $n=30,50$, both RMSE and MAE decreases under $M_{1}$ and $M_{2}$. Overall parametric method performs well in this case. The reason for the poor performance of the semi-parametric method is that the estimates are not accurate, for the parameters associated with the individual degrading which impacts the residual lifetime distributions. Though method $M_{2}$ produce smaller errors, but this methods violates the assumption of our model.
6.4 Case 4
For the case 4, we consider a 3-component mixture of normal distribution for random effect, where mean of first component is 2 and the second component is 4 and third component is 6, where variances of each individual mixture is taken to be small to make each sub-population more homogeneous. Figure 1(d) denotes the degradation observation generated for this case. The mean, standard deviation (s.d) and 95% Credible interval of random effects for each of 10 units, fixed effect $\alpha$ and variance of error $\sigma_{\epsilon}^{2}$ with respect to the different prior combinations for both the parametric and semi-parametric model are given in Tables 19 and 20. The trace plot and autocorrelation plot for the generated mcmc samples of $\alpha$, $\beta_{i}$, $i=1,2,3$ and $\sigma_{\epsilon}^{2}$ for semi-parametric and parametric method are provided in Figure 18 in Appendix B.
Table 19: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$ ), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 2
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 4.738 0.288 (4.181, 5.303) 4.745 0.295 (4.164, 5.331)
$\beta_{2}$ 4.209 0.293 (3.649, 4.802) 4.212 0.307 (3.616, 4.822)
$\beta_{3}$ 2.121 0.294 (1.570, 2.715) 2.131 0.301 (1.566, 2.686)
$\beta_{4}$ 3.411 0.300 (2.808, 3.974) 3.409 0.317 (2.800, 4.002)
$\beta_{5}$ 2.723 0.299 (2.154, 3.303) 2.710 0.315 (2.157, 3.344)
$\beta_{6}$ 5.906 0.304 (5.354, 6.523) 5.920 0.299 (5.362, 6.519)
$\beta_{7}$ 2.534 0.306 (1.922, 3.138) 2.521 0.309 (1.954, 3.142)
$\beta_{8}$ 2.615 0.305 (2.017, 3.232) 2.618 0.307 (2.044, 3.212)
$\beta_{9}$ 5.741 0.307 (5.156, 6.360) 5.742 0.313 (5.152, 6.354)
$\beta_{10}$ 5.863 0.298 (5.298, 6.423) 5.867 0.291 (5.305, 6.436)
$\alpha$ 2.043 0.095 (1.865, 2.221) 2.042 0.095 (1.850, 2.222)
$\sigma_{\epsilon}^{2}$ 0.808 0.091 (0.648, 0.997) 0.810 0.094 (0.649, 1.022)
Table 20: Posterior mean, standard deviation and 95% Credible Interval of random effects ($\beta_{i}$, $i=1,\dots,10$), fixed effect ($\alpha$) and variance of measurement error ($\sigma_{\epsilon}^{2}$) based on Prior 3 for parametric and prior 5 for semi-parametric method
Parameter Parametric method Semi-parametric method
mean s.d 95% Credible interval mean s.d 95% Credible interval
$\beta_{1}$ 4.777 0.301 (4.165, 5.373) 4.774 0.299 (4.104, 5.311)
$\beta_{2}$ 4.205 0.304 (3.600, 4.768) 4.218 0.314 (3.613, 4.848)
$\beta_{3}$ 2.094 0.316 (1.522, 2.720) 2.114 0.302 (1.471, 2.603)
$\beta_{4}$ 3.410 0.317 (2.796, 4.042) 3.404 0.320 (2.843, 3.947)
$\beta_{5}$ 2.715 0.307 (2.132, 3.317) 2.718 0.325 (2.116, 3.264)
$\beta_{6}$ 5.922 0.305 (5.335, 6.535) 5.924 0.280 (5.432, 6.425)
$\beta_{7}$ 2.528 0.300 (1.943, 3.127) 2.526 0.276 (2.070, 3.144)
$\beta_{8}$ 2.609 0.289 (2.023, 3.182) 2.609 0.311 (1.913, 3.132)
$\beta_{9}$ 5.755 0.313 (5.119, 6.343) 5.751 0.313 (5.221, 6.270)
$\beta_{10}$ 5.896 0.300 (5.320, 6.489) 5.885 0.304 (5.344, 6.527)
$\alpha$ 2.040 0.093 (1.856, 2.222) 2.039 0.092 (1.867, 2.205)
$\sigma_{\epsilon}^{2}$ 0.813 0.093 (0.646, 1.005) 0.804 0.089 (0.638, 0.992)
The 95% credible intervals for both proposed semi-parametric and parametric method seems to overestimate true value of error variance, whereas the fixed effect is estimated accurately. The estimates of random effects for both model tends to be similar as number of observations are increased.
Table 21: Estimated median and 95% predictive interval of residual lifetime using a Bayesian parametric method and proposed Bayesian semi-parametric method with $m=31$
Unit True value Parametric Semi-parametric